# Towards improved property prediction of two-dimensional (2D) materials using many-body Quantum Monte Carlo methods

### v1

@@ III.4 Cohesion and adsorption energetics
Since the successful isolation of the graphene sheet, numerous types of 2D materials have been theoretically proposed, each exhibiting intriguing electronic properties. A handful of these materials have also been synthesized. In light of this, obtaining accurate assessments of their ground-state properties, including cohesive energies and their relative differences, is of paramount importance. These assessments not only guide experimental synthesis efforts but also aid in the development of sophisticated DFT exchange-correlation functionals to deal with 2D materials. Quantitative agreement between DMC calculations and experimental results for some existing 2D materials has solidified DMC as the optimal approach for fulfilling this role among current available first-principles calculations for periodic solid systems. DMC calculations have effectively established the ground-state energetics for many proposed 2D materials whose DFT energy differences are sensitive to the choice of density functionals. It is worth noting that, in many instances, DFT tends to overestimate their cohesive energies compared to those derived from DMC calculations [288, 264, 289, 290]. This discrepancy has been attributed to various factors, including dimensional effects and the types of bondings involved. Below, we provide several examples illustrating the application of DMC calculations in investigating the cohesion and adsorption energetics of 2D materials.
@@ III.4.1 Carbon allotropes
Graphyne is one of the proposed sp-sp2 hybridized 2D carbon allotropes expected to have exotic electronic properties, such as possessing a Dirac cone, high carrier mobilities, and a larger surface area than graphene [288]. Since the successful synthesis of graphene, various sp- or sp-sp2 hybridized low-dimensional carbon allotropes have been proposed via various first-principle studies. However, only -graphyne has been experimentally synthesized so far [291, 292, 293], while DFT predicted the possible existence of other stable graphyne structures. To investigate the accurate structural stability of carbon allotropes, QMC is employed on various carbon allotropes, including sp3 diamond, sp2 graphene, sp-sp2 hybridized graphyne, and sp-bonded carbyne chain.
In the results obtained from our DMC calculations for the cohesive energy of carbon allotropes, it is found that the estimated DMC cohesive energies for diamond and graphite are in the excellent agreement with corresponding experimental results [294]. On the other hands, the computed cohesive energy of graphynes from DMC demonstrates a monotonically decreasing trend with the increase of the ratio of sp-bonded atoms on their system. Based on the linear increase of cohesive energy difference between graphene and graphynes with increasing the ratio of sp-bond as shown in Figure 20, we expect that prediction of cohesive energy for other newly proposed graphyne structures can be achieved by utilizing computed DMC bond energies of sp-, sp2-, and sp3- bond as . , and represent cohesive energy, and bond energies for the single bond (), double bound (), and triple bond (), respectively. Note that indicates total number of the carbon atoms (), the single bonds (), the double bonds (), and the triple bonds () per unit cell, respectively. More information regarding this work can be found in Ref. 288.
@@ III.4.2 Boron allotropes
Although no freestanding boron monolayer has been found in nature, a 2D sheet of boron, named borophene, was reported to be synthesized on a metal surface [295, 22, 296, 297, 298]. The structure of borophene is characterized by a hexagonal hole density , the ratio of the number of single-atom vacancies to the total number of atomic sites of the triangular boron sheet. With different values, borophene structures show various combinations of triangular and hexagonal orderings from a complete triangular lattice () to a honeycomb structure (). On the theoretical side, a series of DFT calculations predicted the polymorphism of energetically degenerate borophene structures in the range of and [299, 300]. On the other hand, subsequent experiments have reported the synthesis of borophene phases only outside the polymorphic range predicted by DFT calculations [295, 22, 296, 297, 298], calling for more accurate and systematic investigation of borophene energetics.
Our DMC calculation for bulk -B12 yielded a cohesive energy in very good agreement with its experimental value [301], confirming its accuracy for cohesion energetics of boron allotropes. Subsequent DMC calculations for freestanding borophene structures showed the same polymorphism as predicted by DFT studies (see Fig. 21). Figure 21(a) and (b) present DMC cohesive energies of freestanding borophenes relative to those of -B12 and -sheet borophene (), respectively, along with the corresponding DFT-PBE results. While PBE overestimates energetic stability of borophene against -B12 solid (see Fig. 21(a)), PBE relative energetics between different borophene structures are in quantitative agreement with the DMC results, as seen in Fig. 21(b). This justifies the use of PBE energies to investigate the relative energetics among various borophene structures formed on metal surfaces. The DMC-benchmarked PBE calculations show that as a result of the charge transfer between a metal surface and borophene, the polymorphic range can be expanded to the values of experimentally-synthesized borophenes on top of the Ag(111) and the Au(111) surfaces. Furthermore, the PBE calculations also predicts a possible formation of bilayer borophene with on the Au(111) surface to extend a borophene family. This not only offers further insight into a mechanism of stabilizing 2D borophene structures, but also opens a possibility of borophene-based electronic devices. Additional details of this work can be found in Ref. 290.
@@ III.4.3 Atomic/Molecular adsorption
O2 adsorption on graphene: Capture of oxygen molecules (O2) is important for various industrial applications since O2 can be used to control the rate of combustion, and its rate in the atmosphere should be controlled to avoid corrosion. Graphene has been considered as a suitable substrate for efficient O2 capture, and its adsorption energy was successfully measured experimentally using temperature-programmed terahertz emission microscopy [303]. However, the stable O2 adsorption site and its preferred orientation mode on the surface of graphene has not been confirmed experimentally yet.
@@
Our DMC study of O2 adsorption on graphene found that the O2 orientation mode parallel to the graphene surface is more favorable than the vertical mode. The DMC adsorption energy for two different parallel modes of A and B (see Fig. 22) on the hollow site of the graphene surface is estimated as -0.130(4) eV and -0.126(4) eV, respectively. This similar adsorption energy between A and B indicates possible free planar rotation of O2 at the hollow site. On the other hand, the DMC in-plane diffusion barrier shows that a hollow site is not the most stable adsorption site while vdW-corrected DFT functionals show the lowest adsorption energies at the hollow sites, as seen in Fig. 22. These stable O2 adsorptions at the bridge site of graphene can be understood by the interplay between repulsive interaction and vdW interaction. Our DMC results confirmed that the B orientation mode at the bridge site is the most stable orientation mode for O2 adsorption, and the this mode at the bridge site is hard to accurately describe within the Kohn-Sham framework, even with vdW correction.
@@
Pt clusters on graphene: A graphene-supported Pt cluster has received a great deal of attention because of its enhanced catalytic properties and long-term stability compared to a conventional Pt catalyst [305, 306, 307]. Especially, a single Pt atom catalysis anchored on the graphene surface is considered highly desirable in terms of its reaction efficiency as well as the amount of Pt metal [308, 309]. Nevertheless, its experimental realization has not yet been achieved, and thus it is imperative to understand this system theoretically with thorough investigation of the interaction between graphene and a single Pt atom.
@@
Figure 23 shows our DMC results for the singlet and triplet adsorption energy curves of a single Pt atom adsorbed on three symmetric adsorption sites (a bridge, an on-top, and a hollow site), as a function of the vertical distance from a graphene surface [304]. The bridge-site adsorption with the spin-singlet state turns out to be the most stable, which is consistent with the DFT results. However, it is found that the triplet state becomes energetically preferred over the singlet state for all adsorption sites as the Pt-graphene distance increases, which corresponds to the spin crossing from the chemisorbed singlet state to the physisorbed triplet state. The DMC calculations also predict the presence of local minima in the triplet region. Comparison of the DMC results for the Pt-benzene, the Pt-coronene, and the Pt-graphene system revealed that additive long-range dispersion forces (induced by carbon atoms outside a carbon ring surrounding the Pt atom) are responsible for the formation of the local minima in the adsorption curves. DFT calculations do not capture this local-minimum feature, indicating the significance of many-body correlations at long distances beyond the spin crossing points. This DMC study provides a comprehensive understanding of the Pt adsorption process on a graphene surface.
@@
H2 adsorption on graphenylene: Using hydrogen as a renewable energy resource is a promising route for future technology. Graphenylene, a newly-proposed 2D network of sp2-bonded carbon atoms with large near-circular pores, is a promising membrane for separation of H2 from gas mixtures. Its DMC cohesive energy is estimated to be 6.755(3) eV per atom [310], which is smaller only by meV per atom than the corresponding energy of -graphyne, the most stable structure in a graphyne family. An experimental report of its successful synthesis is understood to reflect this DMC result. DMC calculations are also performed to estimate the adsorption energies of different gas molecules, including H2, on graphenylene, from which the H2 separation capability of a graphenylene membrane against other gas molecules are estimated.
@@
Figure 24(a) presents our DMC adsorption energy of a H2 molecule as a function of the adsorption distance (the vertical distance from a pore center), which was computed for a supercell. The equilibrium adsorption distance is determined through the Morse potential fit of the DMC adsorption energies (see the red dotted line in the figure), which produced the equilibrium distances of 1.97(4) Å for H2. The same procedure results in the DMC equilibrium distances of 2.73(3) Å and 2.73(4) Å for N2 and CO molecules, respectively. After establishing the equilibrium adsorption distances for these molecules, DMC calculations are performed to compute their adsorption energies at both in-plane and equilibrium adsorption distances. Figure 24(b) shows the DMC total energies of the H2-graphenylene complex computed for three different supercell sizes, where the horizontal axis represents the inverse of the number of electrons in a supercell. The adsorption energies of H2 are estimated from the total energies extrapolated to the thermodynamic limit (). The difference between in-plane and equilibrium adsorption energies determines the diffusion barrier of a gas molecule passing through a graphenylene membrane, whose DMC values are estimated to 0.19(2) eV, 0.87(5) eV, and 0.79(2) eV for H2, N2, and CO, respectively. The large difference in diffusion barrier between H2 and other molecules results in extremely high values for hydrogen selectivity against a gas of N2 () or CO () [310]. This suggests that an application of graphenylene as a high performing hydrogen separator is promising.
@@
Atomic H adsoprtion on graphene: The chemisorption of atomic hydrogen on graphene is another system of interest, due to the tunability of electronic (opening up of the band gap [311]) and magnetic (inducing an extended magnetic moment in a graphene sheet [312, 313]) properties when hydrogen atoms are chemisorbed. In addition, graphene and graphitic surfaces can be utilized for hydrogen storage and energy applications [314]. In addition to the interesting applications, the lack of experimental benchmarks (i.e., binding energy) for H chemisorbed on graphene and the variability of standard DFT approaches for atomic adsorption on 2D surfaces makes this an excellent system to apply DMC techniques.
@@
In this work, we performed DFT and DMC calculations to obtain the binding energy of a single H atom chemisorbed on the surface of a graphene sheet [315]. With DMC, we find this binding energy to be -691 meV 19 meV. We find that PBE (plane-wave) overestimates the binding energy by approximately 20 compared to DMC. We also find that PBE0 results in a binding energy close to PBE but HSE yields a binding energy in closer agreement to DMC (-743 meV). We also find significant differences between the DMC and PBE charge densities of graphene and H chemisorbed on graphene (see Fig. 25). More details of this work can be found in Ref. 315. In addition to our work reported in this section that focuses on atomic and molecular adsorption on monolayer surfaces, there have been efforts to utilize DMC to accurately calculate the point defect formation energy in 2D materials [316]. Thomas et al. [316] performed DMC calculations for monovacancies, Stone-Wales defects, and silicon substitutions in monolayer graphene and found substantial disagreements with DFT for defect formation energy (up to 1 eV).
@@ IV Conclusion
We have demonstrated the successes of applying the many-body DMC approach to a wide variety of 2D material systems. This review article provides detailed summaries of several instances where DMC can improve the prediction of magnetic, electronic (including excitonic and topological) properties and accurately capture the interlayer interactions and the energetics of cohesion and adsorption. The results reported in this review demonstrate how accuracy beyond standard DFT can realistically be achieved using many-body electronic structure methods such as DMC. We hope that by showcasing the recent advancements in the field of Quantum Monte Carlo methods being applied to 2D materials, other researchers will be motivated to employ these highly accurate techniques for future work.

### v2

@@ III.4 Cohesion and Adsorption Energetics
Since the successful isolation of the graphene sheet, numerous types of 2D materials have been theoretically proposed, each exhibiting intriguing electronic properties. Several of these materials have also been synthesized. In light of this, obtaining accurate assessments of their ground-state properties, including cohesive energies and their relative differences, is of paramount importance. These assessments not only guide experimental synthesis efforts but also aid in the development of sophisticated DFT XC functionals to deal with 2D materials. Quantitative agreement between DMC calculations and experimental results for some existing 2D materials has solidified DMC as the optimal approach for fulfilling this role among current available first-principles calculations for periodic solid systems. DMC calculations have effectively established the ground-state energetics for many proposed 2D materials whose DFT energy differences are sensitive to the choice of density functionals. It is worth noting that, in many instances, DFT tends to overestimate their cohesive energies compared to those derived from DMC calculations.[349, 314, 350, 351] This discrepancy has been attributed to various factors, including dimensional effects and the types of bondings involved. Below, we provide several examples illustrating the application of DMC calculations in investigating the cohesion and adsorption energetics of 2D materials.
@@ III.4.1 Carbon Allotropes
Graphyne is one of the proposed sp-sp2 hybridized 2D carbon allotropes expected to have exotic electronic properties, such as possessing a Dirac cone, high carrier mobilities, and a larger surface area than graphene.[349] Since the successful synthesis of graphene, various sp- or sp-sp2 hybridized low-dimensional carbon allotropes have been proposed via various first-principle studies. However, only -graphyne has been experimentally synthesized so far,[352, 353, 354] while DFT predicted the possible existence of other stable graphyne structures. To investigate the accurate structural stability of carbon allotropes, QMC is employed on various carbon allotropes, including sp3 diamond, sp2 graphene, sp-sp2 hybridized graphyne, and sp-bonded carbyne chain. We define a parameter , which denotes the ratio of sp-bonded carbon atoms: , where is the number of sp-bonded carbon atoms, and is the total number of carbon atoms in the unit cell. For the carbon allotropes in this study, is 0 for graphene; 0.5, 0.67, and 0.75 for , , and -graphyne, respectively; and 1.0 for carbyne.
In the results obtained from our DMC calculations for the cohesive energy of carbon allotropes, it is found that the estimated DMC cohesive energies for diamond and graphite are in excellent agreement with corresponding experimental results.[355] On the other hand, the computed cohesive energy of graphynes from DMC demonstrates a monotonically decreasing trend with the increase of the ratio of sp-bonded atoms on their system. Based on the linear increase of cohesive energy difference between graphene and graphynes with increasing the ratio of sp-bond as shown in Fig. 20, we expect that prediction of cohesive energy for other newly proposed graphyne structures can be achieved by utilizing computed DMC bond energies of the sp-, sp2-, and sp3- bond as . and represent cohesive energy and bond energies for the single bond (), double bound (), and triple bond (), respectively. Note that indicates the total number of the carbon atoms (), the single bonds (), the double bonds (), and the triple bonds () per unit cell, respectively. More information regarding this work can be found in Ref. 349.
@@ III.4.2 Boron Allotropes
Although no freestanding boron monolayer has been found in nature, a 2D sheet of boron, named borophene, was reported to be synthesized on a metal surface.[356, 22, 357, 358, 359] The structure of borophene is characterized by a hexagonal hole density , the ratio of the number of single-atom vacancies to the total number of atomic sites of the triangular boron sheet. With different values, borophene structures show various combinations of triangular and hexagonal orderings from a complete triangular lattice () to a honeycomb structure (). On the theoretical side, a series of DFT calculations predicted the polymorphism of energetically degenerate borophene structures in the range from to .[360, 361] On the other hand, subsequent experiments have reported the synthesis of borophene phases only outside the polymorphic range predicted by DFT calculations,[356, 22, 357, 358, 359] calling for more accurate and systematic investigation of borophene energetics.
Our DMC calculation for bulk -B12 yielded a cohesive energy in very good agreement with its experimental value,[362] confirming its accuracy for cohesion energetics of boron allotropes. Subsequent DMC calculations for freestanding borophene structures showed the same polymorphism as predicted by DFT studies (see Fig. 21). Fig. 21(a) and (b) present DMC cohesive energies of freestanding borophenes relative to those of -B12 and -sheet borophene (), respectively, along with the corresponding DFT-PBE results. Although PBE overestimates energetic stability of borophene against -B12 solid (see Fig. 21(a)), PBE relative energetics between different borophene structures are in quantitative agreement with the DMC results, as seen in Fig. 21(b). This justifies the use of PBE energies to investigate the relative energetics among various borophene structures formed on metal surfaces. The DMC-benchmarked PBE calculations show that because of the charge transfer between a metal surface and borophene, the polymorphic range can be expanded to the values of experimentally synthesized borophenes on top of the Ag(111) and the Au(111) surfaces. Furthermore, the PBE calculations also predict a possible formation of bilayer borophene with on the Au(111) surface to extend a borophene family. This not only offers further insight into a mechanism of stabilizing 2D borophene structures but also opens a possibility of borophene-based electronic devices. Additional details of this work can be found in Ref. 351.
@@ III.4.3 Atomic/Molecular Adsorption
O2 adsorption on graphene: Capture of oxygen molecules (O2) is important for various industrial applications because O2 can be used to control the rate of combustion, and its rate in the atmosphere should be controlled to avoid corrosion. Graphene has been considered as a suitable substrate for efficient O2 capture, and its adsorption energy was successfully measured experimentally using temperature-programmed terahertz emission microscopy.[363] However, the stable O2 adsorption site and its preferred orientation mode on the surface of graphene have not been confirmed experimentally yet.
@@
Our DMC study of O2 adsorption in triplet-state on graphene [364] found that the O2 orientation mode parallel to the graphene surface is more favorable than the vertical mode. The DMC adsorption energy for two different parallel modes of A and B (see Fig. 22) on the hollow site of the graphene surface is estimated as 0.130(4) eV and 0.126(4) eV, respectively. This similar adsorption energy between A and B indicates possible free planar rotation of O2 at the hollow site. On the other hand, the DMC in-plane diffusion barrier, as seen in Fig. 22, shows that a hollow site is not the most stable adsorption site, while vdW-corrected DFT functionals show the lowest adsorption energies at the hollow sites, as seen in Fig. 22. These stable O2 adsorptions at the bridge site of graphene can be understood by the interplay between repulsive interaction and vdW interaction. Our DMC results confirmed that the B orientation mode at the bridge site is the most stable orientation mode for O2 adsorption, and the mode at the bridge site is hard to accurately describe within the Kohn-Sham framework, even with vdW correction.
@@
Pt clusters on graphene: A graphene-supported Pt cluster has received a great deal of attention because of its enhanced catalytic properties and long-term stability compared to a conventional Pt catalyst.[365, 366, 367] Especially, a single Pt atom catalysis anchored on the graphene surface is considered highly desirable in terms of its reaction efficiency as well as the amount of Pt metal.[368, 369] Nevertheless, its experimental realization has not yet been achieved, and thus it is imperative to understand this system theoretically with thorough investigation of the interaction between graphene and a single Pt atom.
@@
Fig. 23 shows our DMC results for the singlet and triplet adsorption energy curves of a single Pt atom adsorbed on three symmetric adsorption sites (a bridge, an on-top, and a hollow site), as a function of the vertical distance from a graphene surface.[370] Note that the singlet-state many-body wavefunction can be achieved within the single-reference framework of Eq. 2 in the absence of SOC. This is possible because both spin-up and spin-down Slater determinants are constructed from the same Kohn-Sham orbitals obtained from spin-unpolarized DFT calculations. The bridge-site adsorption with the spin-singlet state turns out to be the most stable, which is consistent with our DFT benchmarking results.[370] However, it is found that the triplet state becomes energetically preferred over the singlet state for all adsorption sites as the Pt-graphene distance increases, which corresponds to the spin crossing from the chemisorbed singlet state to the physisorbed triplet state. The DMC calculations also predict the presence of local minima in the triplet region. Comparison of the DMC results for the Pt-benzene, the Pt-coronene, and the Pt-graphene system revealed that additive long-range dispersion forces (induced by carbon atoms outside a carbon ring surrounding the Pt atom) are responsible for the formation of the local minima in the adsorption curves. DFT calculations do not capture this local-minimum feature, indicating the significance of many-body correlations at long distances beyond the spin crossing points. This DMC study provides a comprehensive understanding of the Pt adsorption process on a graphene surface.
@@
H2 adsorption on graphenylene: Using hydrogen as a renewable energy resource is a promising route for future technology. Graphenylene, a newly proposed 2D network of sp2-bonded carbon atoms with large near-circular pores, is a promising membrane for separation of H2 from gas mixtures.[371] Its DMC cohesive energy is estimated to be 6.755(3) eV per atom,[372] which is smaller only by 10 meV per atom than the corresponding energy of -graphyne, the most stable structure in a graphyne family. An experimental report of its successful synthesis [373] is understood to reflect this DMC result. DMC calculations are also performed to estimate the adsorption energies of different gas molecules, including H2, on graphenylene, from which the H2 separation capability of a graphenylene membrane against other gas molecules is estimated.
@@
Fig. 24(a) presents our DMC adsorption energy of a H2 molecule as a function of the adsorption distance (the vertical distance from a pore center), which was computed for a supercell. The equilibrium adsorption distance is determined through the Morse potential fit of the DMC adsorption energies (see the red dotted line in the figure), which produced the equilibrium distances of 1.97(4) Å for H2. The same procedure results in the DMC equilibrium distances of 2.73(3) Å and 2.73(4) Å for N2 and CO molecules, respectively. After establishing the equilibrium adsorption distances for these molecules, DMC calculations are performed to compute their adsorption energies at both in-plane and equilibrium adsorption distances. Fig. 24(b) shows the DMC total energies of the H2-graphenylene complex computed for three different supercell sizes, where the horizontal axis represents the inverse of the number of electrons in a supercell. The adsorption energies of H2 are estimated from the total energies extrapolated to the thermodynamic limit (). The difference between in-plane and equilibrium adsorption energies determines the diffusion barrier of a gas molecule passing through a graphenylene membrane, whose DMC values are estimated to 0.19(2) eV, 0.87(5) eV, and 0.79(2) eV for H2, N2, and CO, respectively. The large difference in the diffusion barrier between H2 and other molecules results in extremely high values for hydrogen selectivity against a gas of N2 () or CO ().[372] This suggests that an application of graphenylene as a high-performing hydrogen separator is promising.
@@
Atomic H adsorption on graphene: The chemisorption of atomic hydrogen on graphene is another system of interest because of the tunability of electronic (opening up of the band gap [374]) and magnetic (inducing an extended magnetic moment in a graphene sheet [375, 376]) properties when hydrogen atoms are chemisorbed. In addition, graphene and graphitic surfaces can be utilized for hydrogen storage and energy applications.[377] In addition to the interesting applications, the lack of experimental benchmarks (i.e., binding energy) for H chemisorbed on graphene and the variability of standard DFT approaches for atomic adsorption on 2D surfaces make this an excellent system to apply DMC techniques.
@@
In this work, we performed DFT and DMC calculations to obtain the binding energy of a single H atom chemisorbed on the surface of a graphene sheet.[378] With DMC, we find this binding energy to be 691 meV 19 meV. We find that PBE (plane-wave) overestimates the binding energy by approximately 20 compared to DMC. We also find that PBE0 results in a binding energy close to PBE, but HSE yields a binding energy in closer agreement to DMC (743 meV). We also find significant differences between the DMC and PBE charge densities of graphene and H chemisorbed on graphene (see Fig. 25). More details of this work can be found in Ref. 378.
Point defect formation energy of graphene: In addition to our work reported in this section that focuses on atomic and molecular adsorption on monolayer surfaces, there have been efforts to utilize DMC to accurately calculate the point defect formation energy in 2D materials.[158] Thomas et al. [158] performed detailed DFT and DMC calculations for point defects in graphene. The types of point defects considered were monovacancies (MVs), silicon substitutions (SiSs), and Stone-Wales (SW) defects. MV involves the removal of one carbon atom, SiS involves replacing a single carbon atom with silicon, and an SW defect is created by rotating a single carbon-carbon bond (in-plane) 90∘ about its midpoint. FS errors of the defect calculations were taken into account by TA and performing the simulations at multiple supercell sizes (ranging from 3 3 to 5 5) and extrapolating to the thermodynamic limit. To reduce the error in TA, the DFT energies were used as a control variate (CV) when computing the DMC energies. A summary of the FS extrapolated results for defect formation energy is depicted in Fig. 26. It was found that DFT (PBE) underestimates the defect formation energy on the order of 1 eV (being slightly smaller for SW and SiS defects). In addition, it was found that the vibrational contribution to the defect formation energy was nonnegligible (on the order of 0.5–1 eV). This work highlights the challenges of accurately computing defect properties with QMC, including systematic and quasirandom finite-concentration effects and significant vibrational contributions. More details of this work can be found in Ref. 158.
@@ IV Conclusion
We have demonstrated the successes of applying the many-body DMC approach to a wide variety of 2D material systems. This review article provides detailed summaries of several instances in which DMC can improve the prediction of magnetic, electronic (including excitonic and topological) properties and accurately capture the interlayer interactions and the energetics of cohesion and adsorption. The results reported in this review demonstrate how accuracy beyond standard DFT can realistically be achieved using many-body electronic structure methods such as DMC. We hope that by showcasing the recent advancements in the field of QMC methods being applied to 2D materials, other researchers will be motivated to employ these highly accurate techniques for future work.